LBDA is a straight angle, Find mLADC 29. (5x+50) (-7x-26)
Answer #1

From the given diagram we can see

ZBDC = (-52 +50°

ZADC = (-7.r - 26)

We know sum of these two angles should be 180o

LADC + LBDC = 180°

Put the values:

(-52 +50° +(-7.x - 26° = 180°

Simplify:

-5.6 + 50 – 7r - 26 = 180

- 12.1 + 24 = 180

-12x=156

r = -13

We need to find LADO.

We know

ZADC = (-7.r - 26)

Put value of x:

ZADC = (-7*(-13) – 26°

ZADC = 91 - 26)

\angle ADC=65^o